OpenStudy (anonymous):

Which numbers should be multiplied to obtain 165^2 - 123^2? a.42 and 42 b.42 and 288 c.1,764 and 1,764 d.4,202 and 4,248

jimthompson5910 (jim_thompson5910):

hint: Difference of squares rule \[\Large a^2 - b^2 = (a-b)(a+b)\]

jimthompson5910 (jim_thompson5910):

165 is not a perfect square but 165^2 is

OpenStudy (anonymous):

is it b? @jim_thompson5910

OpenStudy (anonymous):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

yes you are correct

OpenStudy (anonymous):

can you help me with another one? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure I can do one more

OpenStudy (anonymous):

which of the following is an actual zero of F(x)=x^3+5x^2-4x-20 a.-20 b.-5 c.1 d.4

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I think the answer is b but im not sure.

jimthompson5910 (jim_thompson5910):

a value is a zero (or root) if plugging it in gives 0

jimthompson5910 (jim_thompson5910):

what is f(-20)?

OpenStudy (anonymous):

ohhh so the answer would be A then

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is f(-20)?

jimthompson5910 (jim_thompson5910):

f(x)=x^3+5x^2-4x-20 f(-20)=(-20)^3+5(-20)^2-4(-20)-20 f(-20)=???

OpenStudy (anonymous):

I don't know

jimthompson5910 (jim_thompson5910):

use a calculator to evaluate (-20)^3+5(-20)^2-4(-20)-20

OpenStudy (anonymous):

i got -5940

jimthompson5910 (jim_thompson5910):

if you don't have a calculator, you can type that expression directly into google or wolfram alpha and it will compute it

OpenStudy (anonymous):

i typed it into wolfram alpha and that's what it gave me sooo idk

jimthompson5910 (jim_thompson5910):

good, so because that result is NOT equal to 0, that means x = -20 is NOT a root for f(x)=x^3+5x^2-4x-20

jimthompson5910 (jim_thompson5910):

a value is only a root if you get f(x) = 0

OpenStudy (anonymous):

ok so A is wrong

jimthompson5910 (jim_thompson5910):

yes because f(x) = -5940 when x = -20

OpenStudy (anonymous):

okay what about c?

jimthompson5910 (jim_thompson5910):

what is f(1) ?

OpenStudy (anonymous):

wait I got it the answer is b -5 because (-5)^3+5(-5)^2-4(-5)-20=0 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

thanks for your help :)

jimthompson5910 (jim_thompson5910):

you're welcome